# Why do Machine Learning Notebooks Crash?

### v1

@@ Abstract
Jupyter notebooks have become central in data science, integrating code, text and output in a flexible environment. With the rise of machine learning (ML), notebooks are increasingly used for ML prototyping and data analysis. However, due to their dependence on complex ML libraries and the flexible notebook semantics that allow cells to be run in any order, notebooks are susceptible to software bugs that may lead to program crashes. This paper presents a comprehensive empirical study focused on crashes in ML notebooks. We collect 64,031 ML notebooks containing 92,542 crashes from GitHub and Kaggle, and manually analyze a sample of 746 crashes across various aspects, including exception types and root causes. Our analysis highlights unique root causes related to notebook semantics, including out-of-order execution and previous cell error, that have not been thoroughly covered in earlier research. Furthermore, we categorize crashes as ML bugs or general Python bugs and examine how the crashes are distributed across different stages of the ML pipeline. Our analysis reveals that 87% of crashes are caused by API misuse, data confusion, notebook-specific issues, environment problems, and implementation errors. Crashes are more commonly related to ML bugs than general Python bugs, particularly in Kaggle notebooks, where over 70% of crashes are ML-related. The distribution of crashes across ML pipeline stages differs between the two platforms. Additionally, most crashes (58%) occur during data preparation, model training, and evaluation/prediction stages of the ML pipeline. GitHub and Kaggle exhibit different crash distributions across these stages.
@@
Software bugs, machine learning, Jupyter notebooks, empirical study, crashes
@@ 1 Introduction
Computational notebook environments like Jupyter notebooks have become indispensable in data science due to their interactive nature, offering real-time feedback during code execution [1]. This interactivity makes them ideal for tasks such as data visualization, pattern exploration, and data analysis [2, 3]. Notebooks also streamline documentation by combining text, code, execution results, and images into a single, cohesive environment, designed to enhance collaboration and reproducibility [2, 1].
With the growing availability of data, machine learning (ML) and deep learning (DL) techniques have gained popularity for data modeling and decision-making within the broader field of data science [4, 5]. In this context, ML notebooks have been widely adopted for ML prototyping and data analysis [6]. Hence, ensuring the correctness of ML notebooks becomes essential to prevent software bugs from leaking into the later stages of ML development.
@@
ML notebooks are prone to software bugs with various underlying root causes. Such bugs commonly stem from the heavy reliance on ML libraries that are used for various stages in a ML pipeline including data processing, visualization, and building ML models or other data modeling algorithms [7]. These powerful libraries often come with steep learning curves and varying abstraction levels, leading to errors in API usage, configuration, and compatibility. Furthermore, the flexible semantics of Jupyter notebooks, such as allowing code cells to be executed out of order or multiple times, could increase the likelihood of bugs [2, 1, 8]. Understanding these challenges is crucial for improving the debugging process and developing more robust tools for notebook environments. However, most previous studies [9, 7, 10, 11, 12] have been focused on ML scripts rather than ML programs developed in notebooks.
@@
Bugs in ML programs can manifest as low effectiveness, reduced efficiency, or crashes [9] which abruptly terminate the software. In Python, crashes occur as unhandled exceptions, representing the most severe bug symptom as execution is halted and needs immediate attention. Existing research [7, 12, 8] indicates that crashes are the most common symptoms of bugs in ML programs. Crashes are also more easily measurable due to accompanying error information that can be captured in the notebook environment.
@@
We conduct an empirical study to analyze 92,542 crashes in 64,031 ML notebooks by identifying their exception types, root causes, their association with ML libraries, and the ML pipeline stages where crashes typically occur. Further, we highlight root causes related to the improper use of notebook semantics (e.g., out-of-order execution) which has not been covered in previous research. Our goal is to enhance the use of Jupyter notebooks for ML development, guide the creation of better debugging tools, and improve the overall quality of ML notebooks. To our best knowledge, this is the first study focused on analyzing crashes in ML notebooks.
In particular, we examine crashes in notebooks collected from GitHub and Kaggle, addressing three main research questions. Additionally, we investigate how the characteristics of crashes in ML notebooks differ between these two platforms, providing a crosscutting aspect of our analysis.
@@
RQ1: What are the predominant exception types and root causes of crashes in ML notebooks?
@@
Motivation: Crashes in notebooks often provide detailed error information including exception types that describe the direct trigger of the crash. The root cause, on the other hand, is the underlying reason behind a crash such as wrong usage of APIs, mismatched data expectations, or improper use of notebook semantics (e.g., out-of-order execution). Identifying the most common exception types, root causes, and their relationship help developers and researchers prioritize efforts on handling the most frequent issues in a systematic way. This makes the development of ML applications in notebooks more efficient and productive, while developers can still benefit from the flexibility of Jupyter notebooks.
@@
Results: The most common exception types in ML notebooks are NameError (local or global name not found), ValueError (function receives an argument of correct type but inappropriate value), and TypeError (operation or function is applied to an object with wrong type), attributing to 53% of crashes in total. The main root causes include API misuse (19%), notebook-specific issues (19%), data confusion (18%), environment problems (16%), and implementation errors (15%). Notebook-specific crashes, mostly due to out-of-order execution, are a leading cause on GitHub (20%) and third most common on Kaggle (18%). Our analysis of co-occurrences between exception types and root causes reveal that NameError is primarily (56%) caused by notebook-specific factors. API misuse and data confusion are the primary causes of a wide range of exception types.
@@
RQ2: To what extent are crashes in ML notebooks caused by ML bugs compared with general Python bugs?
@@
Motivation: Crashes in ML notebooks can result from ML-related bugs and general Python bugs. We examine to what extent these crashes are caused by ML bugs, including those directly caused by the wrong use of APIs in ML libraries and other errors that arise from the code using ML APIs or variables. This helps guide better ML API usage practices and improves ML library stability.
@@
Results: ML bugs account for 64% of ML notebook crashes, with 59% on GitHub and a significantly higher proportion (71%) on Kaggle. The primary causes of ML-related crashes are API misuse (23%), data confusion (23%), and notebook-specific issues (21%), while general Python bugs are mostly due to environment (24%) and implementation (24%) errors. API misuse is more common in ML-related crashes compared to Python crashes, highlighting the increased difficulty of using ML libraries correctly.
@@
RQ3: In which stages of the ML pipeline are crashes in notebooks most likely to occur?
@@
Motivation: Studying which stages of the ML pipeline are most prone to crashes, such as data preparation, model construction, or model training stages, helps developers and researchers focus on improving robustness in critical stages, and optimize development resources accordingly.
@@
Results: Crashes predominantly occur during data preparation (27%), model training (16%), and evaluation/prediction (15%) stages. GitHub experiences 3.5 times as many crashes unrelated to any ML pipeline stages compared to Kaggle. The two platforms also exhibit distinct crash distributions across pipeline stages. Additionally, over 70% of the crashes during model training and evaluation/prediction are caused by ML bugs. In data preparation, Kaggle has a higher proportion of ML-related crashes compared to GitHub. In contrast, in the environment setup stage, GitHub notebooks show more ML-related crashes.
Replication package. All artefacts can be found at https://github.com/PELAB-LiU/jupyter_nbs_empirical.
Organization. The remaining of this paper is organized as follows. Section 2 provides background information and discusses related work relevant to our study. Section 3 outlines the experimental setup, detailing the methods used for data collection, processing, and analysis. Next, Section 4 presents the results of our study. Section 5 discusses our main findings and addresses their implications and significance. Finally, Section 6 discusses the potential threats to validity of this study and Section 7 concludes the paper.
@@ 2 Background and related work
Zhang et al. [9] Root cause, symptom ML/DL scripts TensorFlow SO posts, GH issues
Islam et al. [7] Bug type, root cause, symptom, ML pipeline ML/DL scripts Caffe, Keras, TensorFlow, Theano, PyTorch SO posts, GH commits
Humbatova et al. [10] Fault taxonomy (root cause) ML/DL scripts TensorFlow, Keras, PyTorch SO posts, GH commits/issues, interviews, survey
Zhang et al. [11] Failure type, root cause ML/DL scripts Not limited Failure messages from a Microsoft DL platform, interviews
Morovati et al. [12] ML/non-ML bugs, root cause and symptom of ML bugs ML/DL systems TensorFlow, Keras, PyTorch GH issues
@@ 2.1 Terminology and motivating example
First we clarify some key terms following the definitions provided in a previous study [18]. We use bugs or faults as inherent defects that result in incorrect or unexpected behaviors, known as errors. If left unaddressed, these errors can lead to failures or crashes. In Python, crashes are represented by exceptions, typically accompanied by error information.
Jupyter notebooks are the most common medium of ML prototyping and data analysis [6]. Notebooks are more flexible compared to, for example, Python scripts by offering additional interactive features. Figure 1 shows an example excerpt of a ML notebook that trains and evaluates a neural network. A notebook typically consists of multiple cells, which can either contain documentation or executable code. When a user runs a code cell, Jupyter uses a kernel to execute it. After execution, the notebook assigns an execution number to the cell, indicating the order in which it was run. The output (if any) is displayed directly below the code cell. In the example, two code cells have been executed, their outputs are shown below, and the execution numbers appear to the left of each cell, marking the execution order.
@@
The unique flexibility of Jupyter notebooks introduces additional challenges that may affect how crashes occur. Notebooks allow users to execute cells in any order, run the same cell multiple times, and skip executing cells. While this supports iterative and exploratory development, it can lead to unexpected behavior if cells are executed out of sequence. For example, the notebook in Figure 1 can be executed in reverse, such that the code that evaluates the model is executed before the code that trains the model.
@@ 2.2 Studies of bugs in ML or data science programs
Our study addresses these gaps by analyzing crashes in ML notebooks with investigation of how notebook semantics (e.g., out-of-order execution, or previous cell error) impact crashes. Although prior work [2] has linked out-of-order execution to reproducibility issues, we are the first to identify various notebook-specific semantic issues as root causes of crashes. By examining error information in cell outputs, we distinguish ML bugs from general Python bugs and pinpoint the stages of the ML pipeline where these crashes occur.
Most prior research [9, 7, 10, 12, 14, 15, 16, 17, 13, 8] draws from data sources such as Stack Overflow posts, GitHub issues, or code commits. While these sources offer valuable insights, they often describe bugs using natural language, which can abstract away important details about the actual execution context, making it harder to generalize findings. Additionally, Zhang et al. [11] examine DL job failures from an industrial platform. However, job failure messages typically capture issues only after job submission, potentially overlooking critical bugs that occur during coding. In contrast, crash information in Jupyter notebooks provides real-time feedback tied directly to code execution, offering precise and context-rich data. This allows us to capture bugs as they occur, providing more accurate insights into the nature of crashes. By analyzing these crash information, we can identify and diagnose issues more effectively.
Furthermore, many previous studies have focused on bugs in ML/DL programs that use specific ML libraries [9, 7, 10, 13, 12], or defined ML bugs as all errors that occur in ML components in a system [12]. In contrast, our approach is not restricted to a few ML libraries. Instead, we examine a broad range of code from our dataset and identify popular ML libraries. We define ML bugs as bugs in the crashed code that interacts with such libraries, which provides a more direct and precise link between crashes and ML libraries.
@@ 3 Experimental setup
This section outlines the experimental setup of our study, as shown in Figure 2. We begin with the data sources and collection methods, followed by the filtering criteria for identifying the relevant population. Next, we describe the sampling strategy used to select a representative sample. Lastly, we describe the process for manual analysis.
@@ 3.2 Data filtering
This study focuses on notebooks that use Python. To identify the programming language of each notebook, we use the Guesslang library [^1] https://github.com/yoeo/guesslang used in previous works [22, 23] for similar filtering steps. The top three most commonly used languages in GitHub notebooks are Python (95.35%), Julia (1.22%), and R (0.92%). We then filter out non-Python notebooks, leaving a total of 1,020,540 Python GitHub notebooks. All Kaggle notebooks are written in Python.
@@
We aim to study notebook crashes. Hence, we parse and filter all notebooks to retain only those that contain crashes in their cell outputs. Specifically, we filter notebooks containing at least one cell such that 1) its cell_type is code, 2) it comes with output whose output_type is error. After applying this filter, we observe that 10.67% of the collected GitHub notebooks and only 2.61% of Kaggle notebooks contain crashes. This disparity may stem from Kaggle’s role as a platform that encourages developers to share complete, correct, and well-documented notebooks, allowing others to learn from real-world data science challenges and solutions. In total, there are 108,925 GitHub and 4,349 Kaggle notebooks containing 172,638 and 6,745 crashes (i.e., error outputs) respectively.
@@
We exclude notebooks with syntactically incorrect code because such errors prevent the code from running at all, making them irrelevant to the actual crashes developers often face. We focus on crashes that occur during code execution, which reveal deeper issues affecting program quality. To check for syntactic validity, we use Python 3 Abstract Syntax Tree (AST) module, which parses code according to Python 3 grammar. Since Python 3 is the current standard, this approach also filters outdated notebooks, aligning with modern ML workflows.
@@
Jupyter notebooks have various uses. This paper focuses specifically on ML notebooks, defined as those that use at least one ML library. To compile a comprehensive list of ML libraries, we rely on Kaggle as a trusted source, given its reputation as a prominent platform for ML or data science competitions with an active online community. We extract and rank all libraries imported by Kaggle notebooks based on frequency. The top 50 libraries, used by $\sim$95of all Kaggle notebooks, are reviewed by two authors to determine their relevance to ML. From this, we identify 24 popular ML libraries (see Table II), which we use to identify ML notebooks in our dataset.
TABLE II: The 24 most popular ML libraries identified from 166,655 Kaggle notebooks together with their usage frequencies (%), grouped by their main functionality.
@@
Normally, error values of crashes can be extracted from notebooks without extra processing. But, in some cases, developers substitute error values with non-descriptive messages like ignored or null. We replace them with a more descriptive message extracted from the traceback of the crash. When this is not possible, the respective crash is removed from the dataset.

### v2

@@ Abstract
Jupyter notebooks have become central in data science, integrating code, text and output in a flexible environment. With the rise of machine learning (ML), notebooks are increasingly used for prototyping and data analysis. However, due to their dependence on complex ML libraries and the flexible notebook semantics that allow cells to be run in any order, notebooks are susceptible to software bugs that may lead to program crashes. This paper presents a comprehensive empirical study focusing on crashes in publicly available Python ML notebooks. We collect 64,031 notebooks containing 92,542 crashes from GitHub and Kaggle, and manually analyze a sample of 746 crashes across various aspects, including crash types and root causes. Our analysis identifies unique ML-specific crash types, such as tensor shape mismatches and dataset value errors that violate API constraints. Additionally, we highlight unique root causes tied to notebook semantics, including out-of-order execution and residual errors from previous cells, which have been largely overlooked in prior research. Furthermore, we identify the most error-prone ML libraries, and analyze crash distribution across ML pipeline stages. We find that over 40% of crashes stem from API misuse and notebook-specific issues. Crashes frequently occur when using ML libraries like TensorFlow/Keras and Torch. Additionally, over 70% of the crashes occur during data preparation, model training, and evaluation or prediction stages of the ML pipeline, while data visualization errors tend to be unique to ML notebooks.
@@
Software bugs, machine learning, Jupyter notebooks, python, empirical study, crashes
@@ 1 Introduction
Jupyter notebooks, such as those used in Kaggle, JupyterLab, and Google Colab, have become indispensable in data science due to their interactive nature [1]. They offer real-time feedback during code execution, making them ideal for tasks such as data visualization, pattern exploration, and data analysis [2, 3]. Notebooks also streamline documentation by combining text, code, execution results, and visualizations into a single, cohesive environment, designed to enhance collaboration and reproducibility [2, 1].
With the growing availability of data, machine learning (ML) and deep learning (DL) techniques have gained popularity for data modeling and decision-making within the broader field of data science [4, 5]. In this context, Jupyter notebooks, particularly Python-based ones, have been widely adopted for ML prototyping and data analysis [6]. In industrial settings [7], Jupyter notebooks are often the starting point for developing ML applications before transitioning to production-ready ML pipelines, where much of the Python code used in notebooks is reused in large-scale distributed computing environments. This widespread adoption, even in mission-critical applications such as those at NASA Jet Propulsion Lab [8], underscores the importance of quality assurance for notebooks. Bugs that arise during the notebook stage may persist into later development stages, potentially affecting deployed products. Therefore, studying notebook bugs is essential to improve the overall reliability and robustness of ML-driven software.
@@
Software bugs in ML notebooks may be attributed to various underlying root causes. Such bugs commonly stem from the heavy reliance on ML libraries used for various stages in a ML pipeline including data processing and ML model building or other data modeling algorithms [9]. These libraries often come with steep learning curves and varying abstraction levels, leading to errors in API usage, configuration, and compatibility. Furthermore, the flexible semantics of Jupyter notebooks, such as allowing code cells to be executed out of order, could increase the likelihood of bugs [2, 1, 10]. Understanding these challenges is crucial for improving the debugging process and developing more robust tools for notebook environments. However, most previous studies [11, 9, 12, 13, 14] focus on ML scripts, which are standalone, linear code files (e.g., .py files), rather than ML programs developed in notebooks. In this paper, we refer to ML programs as a broader category that includes both ML scripts and ML notebooks.
@@
Bugs in ML programs can manifest in multiple ways, such as low effectiveness, reduced efficiency, or crashes [11] which abruptly terminate the execution. In Python, crashes occur as unhandled exceptions, representing the most severe bug symptom as execution is halted and needs immediate attention. Existing research indicates that crashes are the most common symptoms of bugs in ML programs [9, 14, 10]. Crashes are easier to investigate due to accompanying error information that can be captured in the notebook environment.
@@
We conduct an empirical study to analyze 92,542 crashes in 64,031 publicly available Jupyter notebooks for ML in Python (referred to as ML notebooks throughout the paper). We identify prevalent crash types and root causes, and investigate which ML libraries and pipeline stages most frequently contribute to crashes. Furthermore, we highlight root causes related to the improper use of notebook semantics (e.g., out-of-order execution) which has not been covered in prior research. Our long-term goal is to enhance the use of Jupyter notebooks for ML development, guide the creation of better debugging tools, and improve the overall quality of ML notebooks. To our best knowledge, this is the first study focused on analyzing crashes in ML notebooks.
This study incorporates ML notebooks, used in Jupyter-based environments like Kaggle and JupyterLab, which share the same execution model and .ipynb format. We examine crashes in notebooks collected from representative repositories such as GitHub and Kaggle, which serve distinct user groups with different coding practices. Many data scientists and ML practitioners, particularly in exploratory and collaborative workflows, may lack strong software engineering skills and often commit notebooks with unresolved bugs [10]. By focusing on those notebooks containing crash information, we gain insights into real-world debugging challenges and persistent errors in ML development, which are addressed along three main research questions. We also analyze how crash characteristics vary between GitHub and Kaggle, providing a crosscutting aspect of our analysis.
@@
RQ1: What are the predominant crash types and root causes of crashes in ML notebooks?
@@
Motivation: Crashes in notebooks often provide detailed error information including exception types that describe the direct trigger of the crash. The root cause, on the other hand, is the underlying reason behind a crash such as wrong usage of APIs, mismatched data expectations, or improper use of notebook semantics (e.g., out-of-order execution). Identifying the most common crash types, root causes, and their relationship helps developers and researchers prioritize efforts to handle the most frequent issues in a systematic way. This makes the development of ML applications in notebooks more efficient and productive, while developers can still benefit from the flexibility of Jupyter notebooks.
@@
Results: In our study, the most common crash types in ML notebooks are Variable Not Found (variable name not found), Invalid Argument (function receives an inappropriate argument), and IO Error (errors when handling files), attributing to 37.3% of crashes in total. We identify ML-specific crash types including Tensor Shape Mismatch, Data Value Violation, Out of Memory, Unsupported Broadcast, Feature Name Mismatch, and Model Initialization Error. The main root causes include API misuse (20.9%) and notebook-specific issues (19.4%, mostly due to out-of-order execution). Our analysis reveals that name-related errors (i.e., undefined names) are primarily caused by notebook-specific issues. API misuse, implementation error, and data confusion are the primary causes of a wide range of crash types.
@@
RQ2: To what extent are crashes in ML notebooks caused by ML bugs and by which ML libraries?
@@
Motivation: Crashes in ML notebooks can result from ML-related bugs and general Python bugs. We examine to what extent these crashes are caused by ML bugs, including those that arise from code using ML APIs or involving variables, objects, or other components derived from ML APIs. Additionally, we determine whether a crash results from the direct use of an ML library to identify which libraries are more often associated to crashes and pose greater debugging challenges. This helps guide better ML API usage practices and improves ML library stability.
@@
Results: In our study, ML bugs account for 67.7% of ML notebook crashes, with 62.8% on GitHub and a significantly higher proportion on Kaggle (73.0%). The primary causes of ML-related crashes are API misuse (25%), data confusion (20%), and notebook-specific issues (20%), while general Python bugs are mostly due to implementation (29%) and environment (28%) errors. API misuse is more common in ML-related crashes compared to Python crashes, highlighting the increased difficulty of using ML libraries correctly. In particular, tensorflow/keras and torch are shown to be the most common source of crashes with the highest crash ratio relative to its usage.
@@
RQ3: In which stages of the ML pipeline are crashes in notebooks most frequently to occur?
@@
Motivation: Studying which stages of the ML pipeline are most prone to crashes, such as data preparation, model construction, or model training stages, helps developers and researchers focus on improving robustness in critical stages, and optimize development resources accordingly. This does not imply that all ML notebooks follow a complete pipeline. For example, some may exclusively focus on data preparation, omitting other stages such as model construction or training. Therefore, we analyze which stages of a general ML pipeline caused the crash, regardless of whether a notebook implements all of them.
@@
Results: In our study, crashes predominantly occur during data preparation (33.0%), model training (19.4%), and evaluation/prediction (18.4%) stages. GitHub has more crashes during environment setup and data preparation, whereas more crashes are observed during the evaluation/prediction stage in Kaggle. Additionally, over 70% of the crashes during model training and evaluation/prediction are caused by ML bugs.
Replication package. All experimentation artifacts can be found on GitHub [15], and data on Zenodo [16].
@@ 2 Background and related work
Zhang et al. [11] Root cause, symptom DL scripts TensorFlow SO posts, GH issues
Islam et al. [9] Bug type, root cause, symptom, ML pipeline DL scripts Caffe, Keras, TensorFlow, Theano, PyTorch SO posts, GH commits
Humbatova et al. [12] Fault taxonomy (root cause) DL scripts TensorFlow, Keras, PyTorch SO posts, GH commits/issues, interviews, survey
Zhang et al. [13] Failure type, root cause, ML pipeline DL scripts Not limited Failure messages from a Microsoft DL platform, interviews
Morovati et al. [14] ML/non-ML bugs, root cause and symptom of ML bugs ML(DL) systems TensorFlow, Keras, PyTorch GH issues
@@ 2.1 Terminology and motivating example
First, we clarify some key terms following the definitions provided in a previous study [22]. We use bugs or faults as inherent defects that result in incorrect or unexpected behaviors, known as errors. If left unaddressed, these errors can lead to failures or crashes. In Python, crashes manifest as exceptions, typically accompanied by error information.
Jupyter notebooks are the most common medium of ML prototyping and data analysis [6]. Notebooks are more flexible compared to, for example, Python scripts by offering additional interactive features. Figure 1 shows an example excerpt of an ML notebook that trains and evaluates a neural network. A notebook typically consists of multiple cells, which can contain either documentation or executable code. When a user runs a code cell, Jupyter uses a kernel to execute it. After execution, the notebook assigns an execution number to the cell, indicating the order in which it was run. The output (if any) is displayed directly below the code cell. In the example, two code cells have been executed, their outputs are shown below, and the execution numbers appear to the left of each cell, marking the execution order.
@@
The unique flexibility of Jupyter notebooks introduces additional challenges that may affect how crashes occur. Notebooks allow users to execute cells in any order, run the same cell multiple times, and skip executing cells. While this supports iterative and exploratory development, it can lead to unexpected behavior if cells are executed out of sequence. For example, the notebook in Figure 1 can be executed in reverse, such that the code that evaluates the model is executed before the code that compiles and trains the model. In our example, this would cause a crash.
@@ 2.2 Studies of bugs in ML or data science programs
Our study addresses these gaps by analyzing crashes in Python ML notebooks with investigation of how notebook semantics (e.g., out-of-order execution, or previous cell error) impact crashes. Although prior work [2] has linked out-of-order execution to reproducibility issues, we are the first to identify notebook-specific semantic issues as root causes of crashes. By examining error information in cell outputs, we distinguish ML bugs from general Python bugs and pinpoint the stages of the ML pipeline where these crashes occur.
Most prior research [11, 9, 12, 14, 18, 19, 20, 21, 17, 10] draws from data sources such as Stack Overflow posts, GitHub issues, or code commits. While these sources offer valuable insights, they often describe bugs using natural language, which can abstract away important details about the actual execution context, making it harder to generalize findings. Additionally, Zhang et al. [13] examine DL job failures from an industrial platform. However, job failure messages typically capture issues only after job submission, potentially overlooking critical bugs that occur during coding. In contrast, crash information in Jupyter notebooks provides real-time feedback tied directly to code execution, offering precise and context-rich data. This allows us to capture bugs as they occur, providing more accurate insights into the nature of crashes. By analyzing this crash information, we can identify and diagnose issues more effectively.
Furthermore, many previous studies focused on bugs in ML/DL programs that use specific ML libraries [11, 9, 12, 17, 14], or defined ML bugs as all errors that occur in ML components in a system [14]. In contrast, we examine a broad range of code from our dataset and identify popular ML libraries. As we investigate ML bugs which cause crashes of code that interacts with such libraries, our approach provides a more direct link between crashes and ML libraries.
@@ 3 Experimental setup
This section outlines the experimental setup of our study, as shown in Figure 2. We begin with the data collection methods, followed by the filtering criteria for identifying the relevant population. Next, we describe the sampling strategy used to select a representative sample. Lastly, we describe the process for manual analysis and resampling.
@@ 3.2 Data filtering
This study focuses on Python notebooks. To identify the programming language of each notebook, we use the Guesslang library [26] used in previous works [27, 28] for similar filtering steps. The top three most commonly used languages in GitHub notebooks are Python (95.35%), Julia (1.22%), and R (0.92%). We then filter out non-Python notebooks, leaving a total of 1,020,540 Python GitHub notebooks. All Kaggle notebooks are written in Python.
@@
We aim to study notebook crashes. Hence, we parse and filter all notebooks to retain only those that contain crashes in their cell outputs. Specifically, we filter notebooks containing at least one cell such that 1) its cell_type is code, 2) its output_type is error. After applying this filter, we observe that 10.67% of the collected GitHub notebooks and only 2.61% of Kaggle notebooks contain crashes. This disparity may stem from Kaggle’s role as a platform that encourages developers to share complete, correct, and well-documented notebooks, allowing others to learn from real-world data science challenges and solutions. In total, there are 108,925 GitHub and 4,349 Kaggle notebooks containing 172,638 and 6,745 crashes (i.e., error outputs) respectively.
@@
As our study targets runtime crashes during execution, we exclude notebooks with syntactically incorrect code. Syntax errors prevent execution, making the notebook non-reproducible and unreliable (garbage-in, garbage-out). Furthermore, we investigate notebook-specific issues, such as previous cell errors, focusing on previous cells that execute incorrectly rather than contain syntactically invalid code. Including crashes caused by invalid code would highlight problems that can be resolved using basic static analysis tools, thus yielding no insights into the unique challenges faced during ML notebook development. Additionally, notebooks with persistent syntax errors are typically of lower quality, often containing incomplete work, and may not reflect typical ML development workflows. To check for syntactic validity, we use Python 3 Abstract Syntax Tree (AST) module, which parses code according to Python 3 grammar. Since Python 3 is the current standard, this approach also filters outdated notebooks, aligning with modern ML workflows.
@@
Jupyter notebooks can be used for various applications. This paper focuses specifically on ML notebooks, defined as those that use at least one ML library. To compile a comprehensive list of ML libraries, we rely on Kaggle as a trusted source, given its reputation as a prominent platform for ML or data science competitions with an active online community. We extract and rank all libraries imported by Kaggle notebooks based on frequency. The top 50 libraries, used in $\sim$95of all Kaggle notebooks, are independently reviewed by two authors to determine their relevance to ML. Full agreement is reached in identifying 24 popular ML libraries (see Table II). We then use these libraries to identify ML notebooks in our dataset.
TABLE II: Top 24 popular ML libraries identified from 166,655 Kaggle notebooks with their usage frequencies, grouped by their main functionalities. DL libraries are marked in bold.
@@
Normally, error values of crashes can be extracted from notebooks without extra processing. But, in some cases, developers substitute error values with non-descriptive messages like “ignored” or “null”. We replace them with a more descriptive message extracted from the traceback of the crash. When this is not possible, the respective crash is removed from the dataset.